OpenStudy (anonymous):

and 11-x≥7 what do i do ?

OpenStudy (anonymous):

11-x>=7 -x>=-4 x>=4

OpenStudy (anonymous):

11-x>=7 subtract -11 from both side -x>=-4 divide both side by -1 x<=4

OpenStudy (anonymous):

so what happens to the equals or less then thingy

OpenStudy (anonymous):

we have x<=4 the sign get changed when we divide the function by -1

OpenStudy (anonymous):

oh thanks :

OpenStudy (anonymous):

any other questions of algebra?